OpenStudy (anonymous):

The difference between -7/9 and 4/9 is___? a)-1/3 b)1/3 c)-1 2/9 d)1 2/9

OpenStudy (anonymous):

\[\frac{ -7}{ 9}-\frac{ 4 }{ 9 }= ?\]

OpenStudy (anonymous):

so -7 -4=?

OpenStudy (anonymous):

then put that over 9

OpenStudy (anonymous):

but above is the only choices I have

OpenStudy (anonymous):

simplify afterwards